# Sensor-Based Distributionally Robust Control for Safe Robot Navigation in Dynamic Environments

### v1

@@ 8.3 Sensor-Based CBF Sample Selection
As the robot navigates, we continuously collect range sensor measurements and their corresponding time derivatives, which represent the rate of change of detected distances, denoted as $\frac{\partial\boldsymbol{\eta}}{\partial t}(\phi(\mathbf{x})_{s},\mathbf{R}_{s},t)\in\mathbb{R}^{K}$. This can be estimated by a radar sensor, Doppler LiDAR, or LiDAR velocity algorithms (Yang et al., 2022). In the following, we discuss our approach to obtain samples $\{\boldsymbol{\xi}\}_{i\in[N]}$ by leveraging such measurements.
We collect LiDAR measurements from the last few frames. To account for the robot’s movement, these measurements are transformed to align with the robot’s current pose. From this aggregated data, we select $N$ samples with the lowest values of $\frac{\partial h(\mathbf{x},t)}{\partial t}+\alpha_{h}(h(\mathbf{x},t))$. This criterion is chosen because it effectively identifies those obstacle boundary points where the combined effect of the rate of change in the environment and the current state of the barrier function is most critical. This highlights those samples where the safety constraint, defined in (5), is closest to being violated.
For each selected sample $\{h_{i}(\mathbf{x},t),\frac{\partial h_{i}}{\partial t}\}$, the corresponding gradient $\nabla_{\mathbf{x}}h_{i}(\mathbf{x},t)$ is computed as follows. Let $\mathbf{q}_{i}$ represent the position in the environment corresponding to the LiDAR-detected point for the $i$-th measurement. The vector from the robot’s position $\phi(\mathbf{x})$ to $\phi(\mathbf{x})_{i}$ is given by $\mathbf{v}_{i}=\phi(\mathbf{x})_{i}-\phi(\mathbf{x})$. The gradient of the sample is determined as:
$\nabla_{\mathbf{x}}h_{i}(\mathbf{x},t):=\left[\frac{\mathbf{v}_{i}}{\|\mathbf{v}_{i}\|},0\right]^{\top}.$ (38)
The normalization is essential because the SDF, by its Eikonal property (Crandall and Lions, 1983), has a gradient with a unit 2-norm. The appended zero reflects the fact that the constructed CBF is independent of the robot’s orientation $\theta$. With this procedure, given each $\mathbf{x}\in{\cal X}$, we have available samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ for the vector $\boldsymbol{\xi}$.
@@
With the sample collection strategy for $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ and the CLF design in (29), we formulate the CLF-DR-CBF optimization problem in (18) to synthesize control inputs for the unicycle robot. The unicycle-specific CLF is combined with the DR-CBF constraint, which is based on the CBF samples obtained from the range sensor measurements. This approach enables safe robot navigation in unknown dynamic environments, while directly utilizing the sensor measurements to ensure safety.
@@ 9 Evaluation
We compare our approach with two other safe control strategies, the nominal CLF-CBF QP in (6) and a CLF-Gaussian Process (GP)-CBF second-order cone program (SOCP) (Long et al., 2022). The nominal CLF-CBF QP approach utilizes the closest LiDAR point to define a single CBF $h(\mathbf{x},t)$ and its gradients at each time step. In the CLF-GP-CBF SOCP method, a real-time GP-SDF model (Wu et al., 2021) of the unknown environment is constructed using LiDAR data, from which the CBF, its gradient, and uncertainty information are determined. For a fair evaluation, we solve each of the optimization programs to generate control signals using the Splitting Conic Solver in CVXPY (Agrawal et al., 2018).
In the following simulations and experiments, a consistent set of parameters is utilized to ensure comparability and reproducibility of the results. The linear velocity is constrained between $[-1.2,1.2]$ m/s and the angular velocity is limited within $[-1,1]$ rad/s. The nominal control input $\mathbf{k}(\mathbf{x})$ is set to $[1.2,0]^{\top}$ (the unicycle is commanded to move forward at a speed of 1.2 m/s) and the scaling factor is $\lambda=50$. Table 1 summarizes other parameter values.
Table 1: Simulation and experiment parameters. The class ${\cal K}$ function $\alpha_{V}$ for CLF is assumed to be linear, and $a$ is the linear CBF class ${\cal K}_{\infty}$ function. The parameters $k_{v}$ and $k_{\omega}$ are control gains for linear and angular velocities, respectively, $l$ is the LiDAR offset distance, $\epsilon$ the risk tolerance of the CLF-DR-CBF QP formulation, and $N$ the DR-CBF sample size.
Param. $\alpha_{V}$ $a$ $k_{v}$ $k_{\omega}$ $l$ (m) $\epsilon$ $N$
Value 1.0 0.4 0.05 0.4 0.08 0.1 5
Sec. 9.1 presents simulation results and compares with the two baseline approaches in a static environment in the Gazebo physics simulator (Koenig and Howard, 2004), shown in Fig. 5(a). In Sec. 9.2, we evaluate our approach in dynamic Gazebo environments with pedestrians, shown in Fig. 5(b). Finally, in Sec. 9.3, we test our CLF-DR-CBF QP formulation in cluttered dynamic real-world environments. In all results, the $A^{*}$ algorithm is employed for path planning, operating at a frequency of $5$ Hz. Our CLF-DR-CBF QP formulation is used for real-time obstacle avoidance, running at $50$ Hz.
@@ 9.1 Simulated Static Environments
Figure 6: Comparison of robot trajectories under DR-CBF, GP-CBF, and baseline CBF QP formulations with varying LiDAR noise levels. The planned $A^{*}$ path is shown in green, the actual robot paths in dashed blue, and the robot’s footprint over time as shaded areas. The start and goal points are marked yellow and red, respectively, with an initial orientation of $\theta=0$.
Figure 7: Velocity profile comparisons of different control methods under varying LiDAR noise.
@@
The first set of simulations demonstrates our approach in static environments with varying LiDAR noise, as shown in Fig. 5(a). The robot is tasked to track a planned path (shown in green in Fig. 6) safely in an unknown environment, relying on noisy LiDAR measurements. We introduce Gaussian noise with standard deviation $\sigma$ (ranging from $0.001$ to $0.05$) to each distance measurement of every LiDAR scan. The planned path is not safe for the robot in general, since it does not take the sensing noise into account. Figs. 68 compare our method with the two baselines under different noise levels on the LiDAR measurements. All simulations were conducted on a computer with an Intel i9-12900K CPU with 64 GB RAM.
Table 2: Computation time comparison between different control approaches (in seconds). The values represent the mean $\pm$ standard deviation of the computation time along the trajectory shown in Fig. 6. The total computation time for each method is the sum of the GP map training time (if applicable), inference time, and control synthesis solver time. The CLF-DR-CBF QP and Nominal CLF-CBF QP methods have similar total computation time, as they do not require GP map training. For these two methods, the inference time refers to processing the LiDAR data as CBF samples and corresponding gradients. The CLF-GP-CBF SOCP method has the highest total computation time due to the additional overhead of GP map training.
@@
CLF-DR-CBF QP 0 0.0001 0.0098 ± 0.0027 0.0099 ± 0.0027
CLF-GP-CBF SOCP 0.0086 ± 0.0031 0.0003 0.0119 ± 0.0034 0.0208 ± 0.0065
Nominal CLF-CBF QP 0 0.0001 0.0092 ± 0.0033 0.0093 ± 0.0033
In Figs. 6(a)7(a), and 8(a), we employ our CLF-DR-CBF QP approach to synthesize controls for the robot with Wasserstein radius set to be $r=0.004$. The robot reaches the goal safely in about 14 seconds, demonstrating efficiency and safety in a low-noise regime. Comparing this to the CLF-GP-CBF SOCP strategy (Figs. 6(b)7(b), and 8(b)) under the same noise conditions suggests comparable results in path following and safety. However, as shown in Table 2, our CLF-DR-CBF approach benefits from direct LiDAR data processing, bypassing the GP model’s computational overhead for map training and updates. Besides, as illustrated in Fig. 7(b), during the initial 2-second interval, the robot’s movement under the CLF-GP-CBF SOCP controller is noticeably slower. This reduced speed results from the GP-SDF’s need to accumulate a sufficient volume of LiDAR data for model training, where the initial high estimated CBF variance, due to limited data, constrains rapid movement. Furthermore, as shown in Fig. 8(b), it may be challenging for the GP-SDF to estimate the precise distance to obstacles at around $t=10$s, where the robot is at the top right corner.
When the LiDAR sensor noise level is elevated to $\sigma=0.05$, the variance in the GP-SDF estimation exhibits a substantial increase. This heightened variance can severely undermine the practical applicability of the CLF-GP-CBF SOCP controller, leading to infeasibility of the formulation. Therefore, we do not present the results for the CLF-GP-CBF SOCP controller in this scenario. On the other hand, our CLF-DR-CBF QP controller still works, as shown in Fig. 6(c) to Fig. 8(c). In particular, to account for the larger noise in the measurements, we increased the Wasserstein radius $r$ from $0.004$ to $0.006$. The results underscore the advantage of handling noisy sensor data directly in our formulation without reliance on accurate map estimation.
Comparing Fig. 7(c) to Fig. 7(a), with the CLF-DR-CBF QP controller, we observe that the time needed to complete the task is longer and the control inputs are less smooth due to the larger noise in the LiDAR measurements. However, as shown in Fig. 6(c) and Fig. 8(c), our approach still effectively tracks the planned path and maintains safety. Under the same noise level, we compare our approach with the nominal CLF-CBF QP approach. As shown in Figs. 7(d) and 8(d), the baseline approach requires a longer time to achieve the goal and fails to maintain a safe distance from the obstacle, with a minimum distance of $-0.07$m at around $t=7$s, as evident in Fig. 8(d). This negative distance indicates that the baseline approach violates the safety constraints and comes close to colliding with the obstacle.

### v2

@@ 8.3 Sensor-Based CBF Sample Selection
As the robot navigates through unknown and dynamic environments relying on noisy LiDAR measurements, precisely determining the obstacle set $\mathcal{O}(t)$ and thus computing $h(\mathbf{x},t)$ is not feasible. Therefore, we leverage the distributionally robust CBF formulation from Sec. 6 and obtain samples $\{\boldsymbol{\xi}_{i}\}_{i=1}^{N}$, where recall
$\boldsymbol{\xi}=[\mathbf{F}^{\top}(\mathbf{x})\nabla_{\mathbf{x}}h(\mathbf{x},t),\;\alpha_{h}(h(\mathbf{x},t)),\;\frac{\partial h(\mathbf{x},t)}{\partial t}],$
directly from the distance measurements $\boldsymbol{\eta}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x}))=[\eta_{1}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x})),\dots,\eta_{K}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x}))]^{\top}$. Note that the system dynamics $\mathbf{F}(\mathbf{x})$ and ${\cal K}_{\infty}$ function $\alpha_{h}$ are assumed to be known and deterministic.
In practical robotic systems, sensor measurements, such as LiDAR readings, are collected at discrete intervals determined by the sensor’s frequency. These measurements must be processed, along with state estimation and control synthesis, within the time constraints of the control loop to ensure real-time applicability.
To account for state estimation uncertainty, we consider $M$ samples of the robot’s estimated pose, denoted as $\{\mathbf{x}^{(j)}\}_{j=1}^{M}$, where each sample represents a possible true pose of the differential-drive robot given its localization error distribution. For each sample $\mathbf{x}^{(j)}$, we have $K$ corresponding LiDAR measurements in the robot’s local frame, $\{\boldsymbol{\eta}_{i}^{(j)}\}_{i=1}^{K}$. These measurements are transformed into the global frame using the estimated pose, yielding LiDAR hit $\mathbf{q}_{i}^{(j)}(t)\in\mathbb{R}^{2}$ with $i$-th LiDAR measurement of pose $\mathbf{x}^{j}$ at time $t$. By aggregating these transformed measurements, we construct a comprehensive set of LiDAR points in the global frame:
$\mathcal{P}(t)=\left\{\mathbf{q}_{i}^{(j)}(t)\mid i=1,\dots,K;\;j=1,\dots,N\right\}\subset\mathbb{R}^{2},$ (46)
resulting in a total of $N\times K$ points. This set effectively captures the combined uncertainties in both state estimation (localization) and sensor measurements. To account for dynamic obstacles, we may obtain estimates of the time derivatives $\{\frac{\partial\mathbf{q}_{i}^{j}}{\partial t}\}$ using a radar sensor, Doppler LiDAR, or a LiDAR velocity estimation algorithm (Yang et al., 2022).
Since we cannot compute $h(\mathbf{x},t)$ exactly, we use the available LiDAR points set ${\cal P}(t)$ to approximate the barrier function and its gradients. From this aggregated data, we select $N$ samples that minimize the following criterion:
$\frac{\partial}{\partial t}$ $d_{\mathcal{B}_{0}}\left(\mathbf{R}(\mathbf{x})^{\top}\left(\mathbf{q}_{i}(t)-\phi(\mathbf{x})\right)\right)+$
$\alpha_{h}\left(d_{\mathcal{B}_{0}}\left(\mathbf{R}(\mathbf{x})^{\top}\left(\mathbf{q}_{i}(t)-\phi(\mathbf{x})\right)\right)\right),$ (47)
for $i=1,\dots,K$. Note that the generalized gradient $\partial h_{i}(\mathbf{x},t)$ can be computed correspondingly by utilizing the robot SDF $d_{{\cal B}_{0}}$. This criterion effectively identifies obstacle points where the combined effect of the barrier function’s rate of change and its current value is most critical, highlighting the samples where the safety constraint, as defined in (5), is closest to being violated.
With this procedure, given each $\mathbf{x}\in{\cal X}$, we have available samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ for the vector $\boldsymbol{\xi}$.
@@
With the sample collection strategy for $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ and the CLF design in (33), we formulate the CLF-DR-CBF optimization problem in (21) to synthesize control inputs for the unicycle robot.
The differential-drive specific CLF is combined with the DR-CBF constraint, which is based on the CBF samples obtained from the range sensor measurements and estimated robot poses. This approach enables safe robot navigation in unknown dynamic environments, while directly utilizing the sensor measurements to ensure safety.
@@ 9 Evaluation
We compare our approach with two other safe control strategies, the nominal CLF-CBF QP in (6) and a CLF-Gaussian Process (GP)-CBF second-order cone program (SOCP) (Long et al., 2022). The nominal CLF-CBF QP approach utilizes the closest LiDAR point to define a single CBF $h(\mathbf{x},t)$ and its gradients at each time step. In the CLF-GP-CBF SOCP method, a real-time GP-SDF model (Wu et al., 2021) of the unknown environment is constructed using LiDAR data, from which the CBF, its gradient, and uncertainty information are determined. While the GP-SDF mapping process contributes to safety by continuously updating the environment representation, it also incurs computational overhead due to the real-time update of the GP-SDF model. For a fair evaluation, we solve each of the optimization programs to generate control signals using the Interior Point Optimizer through the CasADi framework (Andersson et al., 2019).
In the following simulations and experiments, a consistent set of parameters is utilized to ensure comparability and reproducibility of the results. The linear velocity is constrained between $[-1.2,1.2]$ m/s and the angular velocity is limited within $[-1,1]$ rad/s. The nominal control input $\mathbf{k}(\mathbf{x})$ is set to $[1.2,0]^{\top}$, directing the robot to move forward at $1.2$ m/s. While we use a constant nominal controller, note that our formulation supports more complex, state-dependent nominal controllers. The scaling factor is $\lambda=50$. Table 1 summarizes other parameter values.
Table 1: Simulation and experiment parameters. The class ${\cal K}$ function $\alpha_{V}$ for CLF and the class ${\cal K}_{\infty}$ for CBF are assumed to be linear. The parameters $k_{v}$ and $k_{\omega}$ are control gains for linear and angular velocities, respectively, $\epsilon$ the risk tolerance of the CLF-DR-CBF QP formulation, and $N$ the DR-CBF sample size.
Parameters $\alpha_{V}$ $\alpha_{h}$ $k_{v}$ $k_{\omega}$ $\epsilon$ $N$
Value 1.0 1.5 0.05 0.4 0.1 5
The layout of this section is as follows. Sec. 9.1 presents simulation results and compares with the two baseline approaches in static environments (e.g., Fig. 5(a)). In Sec. 9.2, we evaluate our approach in dynamic Gazebo environments (Koenig and Howard, 2004) with pedestrians, shown in Fig. 5(b). Finally, in Sect. 9.3, we test our CLF-DR-CBF QP formulation in dynamic cluttered real-world environments. In all results, the $A^{*}$ algorithm is employed for path planning, operating at a frequency of $5$ Hz. Our CLF-DR-CBF QP formulation is used for real-time safe navigation, running at $50$ Hz.
@@ 9.1 Simulated Static Environments
Figure 6: Evaluation of robot navigation performance under varying conditions. The simulations demonstrate the system’s behavior across different robot shapes and noise conditions, with (a) showing the tested robot shapes, (b) illustrating the impact of shape variation on navigation, and (c-d) analyzing the effects of sensor and localization uncertainties on trajectory execution. For (c) and (d), Shape 1 from (a) is used as the test case.
Table 2: Performance metrics for static environments under varying LiDAR and localization noise. The metrics include stuck rate and collision rate out of 1000 trials and tracking error (m) (mean $\pm$ std). The stuck rate reflects cases of infeasibility or the robot being trapped in local optima. The robot starts from the origin, and goals are placed at least 10 meters away in Figure 5(a).
LiDAR Noise Localization Noise Method Stuck Rate Collision Rate Tracking Error (m)
$\sigma=0.001$ $\sigma=0.01$ CLF-DR-CBF QP 0.0 0.0 1.33 $\pm$ 0.19
CLF-GP-CBF SOCP 0.2 0.0 1.38 $\pm$ 0.31
Nominal CLF-CBF QP 0.0 0.2 1.26 $\pm$ 0.20
$\sigma=0.05$ CLF-DR-CBF QP 0.1 0.0 1.88 $\pm$ 0.35
CLF-GP-CBF SOCP 0.3 15.8 2.16 $\pm$ 0.47
Nominal CLF-CBF QP 0.1 15.5 1.92 $\pm$ 0.34
$\sigma=0.05$ $\sigma=0.01$ CLF-DR-CBF QP 0.0 0.0 1.77 $\pm$ 0.38
CLF-GP-CBF SOCP 25.1 0.2 1.97 $\pm$ 0.52
Nominal CLF-CBF QP 1.7 4.4 1.82 $\pm$ 0.41
$\sigma=0.05$ CLF-DR-CBF QP 1.8 0.7 2.22 $\pm$ 0.57
CLF-GP-CBF SOCP 31.3 13.9 2.58 $\pm$ 0.78
Nominal CLF-CBF QP 3.3 25.8 2.23 $\pm$ 0.59
$\sigma=0.1$ $\sigma=0.01$ CLF-DR-CBF QP 1.2 0.0 2.31 $\pm$ 0.66
CLF-GP-CBF SOCP 60.7 0.0 2.55 $\pm$ 0.79
Nominal CLF-CBF QP 4.7 7.7 2.52 $\pm$ 0.72
$\sigma=0.05$ CLF-DR-CBF QP 6.6 1.8 2.61 $\pm$ 0.75
CLF-GP-CBF SOCP 65.4 22.0 2.78 $\pm$ 0.91
Nominal CLF-CBF QP 8.9 58.3 2.55 $\pm$ 0.69
The first set of simulations aims to validate the robustness and adaptability of the proposed CLF-DR-CBF QP formulation in static environments.
Hypothesis: Our CLF-DR-CBF QP method can ensure safe navigation by dynamically adjusting to varying robot shapes and compensating for sensor and localization noise.
Setup: The robot is tasked to follow the planned path while avoiding obstacles. Gaussian noise with varying standard deviations ($\sigma$) is added to the LiDAR measurements, while the localization error is modeled as a Gaussian random vector with varying standard deviation levels as well. Diverse robot shape geometries (Figure 6(a)) are considered to evaluate the method’s adaptability. Unless otherwise noted, the default robot shape is shape 1, representing the original Jackal robot. The trajectories are analyzed under different conditions to assess safety, adaptability, and smooth navigation.
Results and discussion: Figure 6(b) demonstrates the trajectories for each robot shape. The results show that our approach dynamically adjusts the robot’s path based on its geometry, ensuring safety by maintaining adequate clearance from obstacles while achieving smooth navigation. For symmetric shapes like robot shape 1 (original Jackal robot), the executed path is closer to the obstacle boundary. In contrast, for highly asymmetric shapes like robot shape 3, the executed path deviates more significantly to account for the robot’s geometry to ensure safety. Unless otherwise noted, the simulations presented in the following are performed using robot shape 1.
Figures 6(c) and 6(d) explore the effects of sensor and localization noise on navigation performance. In Figure 6(c), our CLF-DR-CBF QP method successfully compensates for sensor inaccuracies, maintaining safety even at higher noise levels ($\sigma=0.2$). Similarly, Figure 6(d) evaluates the impact of localization noise on trajectory execution, where the robot effectively adapts to inaccuracies in state estimation, showcasing the robustness of our distributionally robust formulation.
We next evaluate the performance of our CLF-DR-CBF QP formulation in static environments simulated in Gazebo (Fig. 5(a)), comparing it with two baseline approaches: Nominal CLF-CBF QP and CLF-GP-CBF SOCP (Long et al., 2022).
Hypothesis: Our CLF-DR-CBF QP formulation ensures robust and safe navigation under varying levels of sensor and localization noise. Compared to baseline methods, our approach should achieve lower failure rates (stuck and collision) and demonstrate superior adaptability to noise conditions without compromising computational efficiency.
Setup: The simulations are conducted in a static Gazebo environment where the robot is tasked to achieve a predefined goal while avoiding obstacles. Gaussian noise with standard deviation $\sigma$ (ranging from 0.001 to 0.1) is added to the LiDAR measurements, and localization noise with $\sigma$ values up to 0.05 is introduced. For each noise level, 1000 trials are conducted with randomly placed goal locations at least 10 meters away from the robot’s starting position. The evaluation metrics include stuck rate, collision rate, and average tracking error (mean $\pm$ std), as summarized in Table 2.
The stuck rate and collision rate together determine the success rate, representing the percentage of trials where the robot successfully reaches the goal without safety violations. The stuck rate captures two failure modes. First, the optimization program may become infeasible due to large uncertainties, particularly in the GP-CBF method, where the GP-SDF map’s high variance at large LiDAR noise levels often renders the SOCP problem infeasible. Second, the robot may become trapped in a corner (local minimum), unable to make further progress toward the goal.
@@
Table 3: Computation time comparison between different control approaches (in seconds). The values represent the mean $\pm$ standard deviation of the computation time along the robot trajectory. The total computation time for each method is the sum of the GP map training time (if applicable), inference time, and control synthesis solver time. The CLF-DR-CBF QP and Nominal CLF-CBF QP methods have similar total computation time, as they do not require map updates. For these two methods, the inference time refers to processing the LiDAR data as CBF samples and corresponding gradients. The CLF-GP-CBF SOCP method has the highest total computation time due to the additional overhead of GP map training.
@@
CLF-DR-CBF QP 0 0.0002 0.0071 $\pm$ 0.0022 0.0073 $\pm$ 0.0022
CLF-GP-CBF SOCP 0.0086 $\pm$ 0.0031 0.0003 0.0096 $\pm$ 0.0028 0.0185 $\pm$ 0.0059
Nominal CLF-CBF QP 0 0.0002 0.0064 $\pm$ 0.0023 0.0066 $\pm$ 0.0023
Results and discussion: The results in Table 2 highlight the robustness of the CLF-DR-CBF QP method, which consistently achieves low stuck and collision rates, even under high noise levels. In contrast, the GP-CBF and Nominal CLF-CBF QP methods exhibit significant performance degradation in challenging noise conditions.
As localization noise increases, the two baseline methods are more prone to collisions. This is because their formulations do not explicitly account for uncertainties in the robot state estimation. Our CLF-DR-CBF QP method remains robust by explicitly addressing localization uncertainties in its formulation. Similarly, at higher LiDAR noise levels, the GP-CBF method struggles due to significant variance in the GP-SDF estimation, often rendering its optimization program infeasible. The CLF-DR-CBF QP method, by directly using LiDAR measurements and incorporating distributionally robust constraints, avoids reliance on explicit map construction, exhibiting a higher probability of reliable performance at higher noise levels.
Figure 7 complements these quantitative findings by reporting the estimated distance to obstacles, true distance, and path tracking errors. Under low noise conditions ($\sigma=0.001$), the CLF-DR-CBF QP approach achieves smooth and safe navigation, successfully tracking the planned path and maintaining safety. The GP-CBF method performs similarly in this scenario but incurs higher computational overhead due to the GP map training and inference, as shown in Table 3.
As the noise level increases to $\sigma=0.05$, the GP-CBF method becomes infeasible (stuck) due to significant variance in the GP-SDF estimation. In this scenario, the CLF-DR-CBF QP formulation adapts by increasing the Wasserstein radius $r$ from $0.004$ to $0.006$, enabling it to handle the noisier sensor data effectively. Although higher noise results in slightly longer task completion times, the CLF-DR-CBF QP method ensures safety and successfully completes the task. The Nominal CLF-CBF QP approach, however, fails to maintain safety, with recorded minimum distances dropping below zero in Figure 7(d) (e.g., $-0.07$m at $t=7$s).
Table 3 highlights the computational efficiency of the proposed method. The CLF-DR-CBF QP formulation achieves computation times comparable to the Nominal CLF-CBF QP method while significantly outperforming the GP-CBF method. This advantage stems from the CLF-DR-CBF QP method’s direct use of LiDAR measurements without requiring computationally expensive GP map construction.
Overall, these results demonstrate the robustness and efficiency of the proposed CLF-DR-CBF QP formulation. By directly handling noisy sensor data and avoiding reliance on explicit map reconstruction, the method effectively balances computational efficiency and robust safety, making it suitable for real-world applications where sensing noise, localization noise, and computational constraints are significant challenges.